# Cluster Synchronization via Graph Laplacian Eigenvectors

### v1

@@ Abstract
Almost equitable partitions (AEPs) have been linked to cluster synchronization in oscillatory systems, providing a mathematical framework for understanding collective behavior. Using the spectral properties of AEPs, we can describe this synchronization behavior in terms of the graph Laplacian eigenvectors. Our results also shed light on transient hierarchical clustering as well as multi frequency clustering, and the conditions under which they can occur. Through our analysis we are able to relate dynamical clustering phenomena directly to network symmetry and community detection, joining the structural and functional notions of clustering in complex networks. This bridges a crucial gap between static network topology and emergent dynamic behavior. Additionally, this description of the problem allows us to define a relaxation of an AEP we call a quasi-equitable partition (QEP). Perfect AEPs are rare in real-world networks since most have some degree of irregularity or noise. While relaxing these strict conditions, QEPs are able to maintain many of the useful properties allowing for qualitatively similar clustering behavior as with AEPs. Our findings have important implications for understanding synchronization patterns in real-world networks, from neural circuits to power grids.
@@ I Introduction
The study of complex networks and their dynamical behaviors has been a fertile ground for understanding synchronization phenomena across different fields. From meso-scale neural activation, to power flow networks, numerous complex networks exhibit interacting oscillatory behavior. In the study of these systems, it is often natural to ask what sort of synchronization behaviors can occur. In recent years, there has been evidence from neuroscience of an unexplored relationship between resting state clusters in the brain and the Laplacian eigenvectors of functional connectivity networks [1]. Existing theory does not sufficiently explain the correspondence between functional clustering and structurally defined eigenvectors, particularly in the presence of high amounts of noise. Our work approaches this question from the perspective of network oscillatory models, closing this gap between structural and functional clusters.
Pioneering work by Kuramoto [2] on what is now called the Kuramoto model, brought fundamental insights into how coupled oscillators spontaneously synchronize, and laid the groundwork for subsequent investigations into network-level synchronization dynamics [3]. The Kuramoto model and its derivatives are used as an abstraction of oscillatory behavior for study. This model represents vertices as oscillators and couples their differential equations over the edges of the network. Consider a weighted graph $G$ with adjacency matrix $A$. For every vertex $i$, define the natural frequency as $\omega_{i}\in\mathbb{R}$. Then the phase of vertex $i$ over time is determined by the differential equation
@@
Naturally, the synchronization characteristics of a Kuramoto model are informed by the structure of the graph it is placed upon. However, the nature of this relationship is unclear, particularly due to the multitudinous structures a graph can take. Recent advances in the study of coupled oscillator dynamics have progressively refined our understanding of how graph structure influences synchronization potential [4] and have proven that Almost Equitable Partitions (AEPs) facilitate cluster synchronization via their quotient graph. Specifically, the characterization of synchronized clusters in networked oscillators has highlighted the intricate relationship between graph structure and dynamical coherence [5]. Simultaneously, in more general signal processing literature, a novel approach has been emerging, which focuses on the discretization of the Fourier analysis framework to better study the fundamental tradeoff between a signal’s localization to the vertices of a graph and in its spectral domain [6].
@@
Despite these significant advances, a crucial gap persists between the spectral properties of AEPs, studied in graph theory and network science, and the cluster synchronization they allow in complex networks. Traditionally, their spectral properties have proven to be crucial for the detectability of AEPs on graphs [7], but have till this point found little purchase in the space of vertex localized signals on graphs and importantly for us, the dynamics of cluster synchronization. While extensive research has explored fully equitable partitions, the behavior of systems that nearly -but not quite— satisfy these structural conditions remains largely unexplored. Existing literature, such as works by Arenas [8] and Breakspear [9], have hinted at the complexity of synchronization in networks with subtle structural variations, but a comprehensive spectral framework for analyzing these near-equitable configurations has been lacking.
@@
Building upon spectral graph theory and drawing from recent developments in network synchronization research, we propose a generalized framework that extends traditional equitable partition analysis. Researchers like Schaub, and Thibeault [10, 11, 12] have provided foundational insights into matrix reduction and system dimensionality that inform our approach. By leveraging eigenvector interpretation techniques [6], we develop a more nuanced understanding of how network structures facilitate or inhibit synchronization.
@@
In the sections that follow, we build the necessary foundation for an understanding of cluster synchronization in the Kuramoto model, entirely through the information encoded in its graph Laplacian. We begin by considering the structural implications within the Laplacian’s spectrum, and particularly the role of partitions therein (Section 2). We then discuss the relationship between the spectrum of the Laplacian and its dynamics. Following this background, we present a novel spectrally motivated description of cluster synchronization, and discuss its implications to the study of stability for weighted graphs (Section 3). The presented spectral description of clustered dynamics allows for new interesting applications. We look into two examples: multi-frequency clustering, and modeling with antisymmetric phase lag. Lastly, Section 5 offers concluding remarks and future research directions.
@@ II The Spectral description of Synchronized Clusters
The field of spectral graph theory involves itself with the linear algebra of graph representations and the relationship with the graph structure. This crucial foundation for the field of network analysis allows us to leverage matrix algebra tools, built for computing, to understand graphs at scales we cannot feasibly study combinatorially.
Matrix representations of graphs come in many forms, and we present a few with relevance to our application. Let $G=(V,E,w)$ be a weighted connected graph with $n$ vertices $v_{1},\dots,v_{n}\in V$ and $m=|E|$ edges, where $w:E\to\mathbb{R}^{+}$ is the weight function for the edges. Perhaps the most natural description of this graph comes in the form of the adjacency matrix $A\in\mathbb{R}^{n\times n}$, whose $ij$th entry is $w((v_{i},v_{j}))$ if $(v_{i},v_{j})\in E$ and 00 otherwise.
Another important representation comes in the form of the signed incidence matrix, also known as the boundary matrix. This matrix ignores weight assignments, and requires an arbitrary orientation on the edges of our graph. For simplicity, we induce an orientation from the vertex order. Define the signed incidence matrix as $B\in\mathbb{R}^{n\times m}$ such that $b_{ie}=1$ if edge $e=(i,j)$ and $i<j$, and $b_{ie}=-1$ if $i>j$. Finally, $b_{ie}=0$ if edge $e$ is not incident to vertex $i$. This boundary matrix serves as a fundamental building block of the spectral form of the Kuramoto equation shown later. We will see that when viewed as an operator, it offers a bridge between vertex values and edge values.
Finally we introduce the matrix we pay particular attention to in this paper, the graph Laplacian. We define the graph Laplacian as $L=D-A$, where $A$ is the adjacency matrix and $D\in\mathbb{R}^{n\times n}$ is the diagonal matrix of the row sums of $A$. The entries of $D$ are the sums of incident edge weights. In the case of uniform weight $1$ (an ’unweighted’ graph), they equal the degrees of the corresponding vertices. The graph Laplacian is a cornerstone in spectral graph theory and has numerous applications in network analysis, clustering, and graph-based dynamical systems [13].
@@
It is well-established that the eigenvectors of the Laplacian play a crucial role in revealing organizational structure in graphs, particularly in terms of their symmetry and connectivity patterns. For example, the multiplicity of the zeroth eigenvalue reveals the number of connected components and in, the case of multiplicity $1$, its corresponding normalized eigenvector is $v_{0}=\frac{1}{\sqrt{N}}\mathbb{1}$.
@@ II.1 Spectrum and Almost Equitable Partitions
One frame under which to understand symmetries of graphs or groupings of structurally similar vertices, is almost equitable partitions (AEP) of a graph, also known as an externally equitable partition.
We define an AEP on an unweighted graph before proceeding to our weighted case. Let $N(v)$ denote the set of neighbors of $v\in V$. An AEP is a partition of the vertex set $\pi=\{V_{1},\dots,V_{k}\}$ such that for $V_{i},V_{j}\in\pi$ where $i\neq j$, there exists $d_{ij}\in\mathbb{N}$ such that for all $v\in V_{i}$, we have that $|N(v)\cap V_{j}|=d_{ij}$ (See fig.1 B.1). In other words, the partition cell a vertex is in determines the number of neighbors that vertex has in each other partition cell. Such a partition encodes a grouping of vertices that play the same ‘role’ in the connectivity of the graph between the delineated partition cells. Note that the almost equitable partition ignores any connectivity within a partition cell, the case in which this too is regular is called an equitable partition (EP). The lack of constraints within partitions cells of an AEP allow for them to exist at multiple, overlapping, scales.
Now we generalize this notion and define an AEP in the context of our weighted graph [5]. Let $G=(V,E,w)$ be a weighted graph with weight function $w:E\to\mathbb{R}^{+}$. Then a weighted almost equitable partition of $G$ is a vertex partition $\pi=\{V_{1},\dots,V_{k}\}$ such that for $V_{i},V_{j}\in\pi$ with $i\neq j$, there exists $d_{ij}\in\mathbb{R}^{+}$ such that for all $v\in V_{i}$ we have $\sum_{w\in N(v)}w((v,w))=d_{ij}$. In other words, a weighted AEP is again a vertex partition with regular inter cell connectivity, except now the regularity comes in the form of a constant sum of edge weights rather than a constant number of neighbors. Note that we recover the original definition of an AEP with an unweighted graph. For the rest of this paper our discussions on AEPs refer to the weighted AEP unless otherwise specified.
We can use this notion of vertex partitioning, or vertex role assignment, to coarse-grain the graph, summarizing the implied structure in what is known as the quotient graph (See fig. 1 C.1). Consider a graph $G$ with arbitrary partitioning $\pi$. Define the indicator matrix of $\pi$ as $P^{\pi}\in\mathbb{R}^{n\times k}$ such that $P^{\pi}_{il}=1$ if vertex $i$ is in partition cell $V_{l}$, and 00 otherwise. Note we omit the $\pi$ superscript when the choice of partition is clear. The rows of $P$ correspond to the vertices of $G$ and its columns correspond to the partition cells of $\pi$. Now given an $n$ by $n$ matrix representation of a graph like the graph Laplacian (or adjacency), $L$, we define the quotient Laplacian (or adjacency) matrix of $L$ with respect to the partition $\pi$ as
@@
Figure 1: In A.1 we see a graph with the associated graph laplacian A.2. One may observe an AEP into three parts on the graph, color coded in B.1. The qoutient graph associated with this AEP is shown in C.1, with the eigenvectors of the quotient laplacian seen in C.2. Observing the eigenvectors of the graph laplacian in B.2, we see the scaled structural eigenvectors from C.2 on the left hand side.
The quotient graph provides a powerful tool for coarsely analyzing the structure of the original graph, while preserving connectivity information relevant to the partition. Indeed, cluster synchronization can be understood as a reduction of the dynamics onto such a quotient graph [4]. The preservation of information is made clear in the analysis of the quotient Laplacian in the context of an almost equitable partition. Notice, in the case of an almost equitable partition $\pi$, the off diagonal blocks of the induced block partition of $L$ have average row sums exactly equal to one another. This implies that the off diagonal entries of $L^{\pi}$, which correspond to the average of these row sums, preserve all the information of inter cell connectivity. The following theorem states this algebraically.
@@
[14, 15, 16] Let $L$ be the Laplacian of a graph $G=(V,E,w)$, and $\pi=\{C_{1},\cdots,C_{k}\}$ be a partition of $V$. Then $\pi$ is an almost equitable partition if and only if the image of $P^{\pi}$ is $L$ invariant. In this case with $L^{\pi}=(P^{T}P)^{-1}P^{T}LP$ we have $LP=PL^{\pi}$.
Theorem 1 comes from a more general theory of the tight interlacing the eigenvalues of $L^{\pi}$ have with respect to the eigenvalues of $L$ [14]. Notice that given, $(\lambda,v)$, an eigen-pair of $L^{\pi}$, it follows
@@
[14, 16] Let $\pi$ be an almost equitable partition of a graph, and $P$ the indicator matrix associated to $\pi$. Then $(v,\lambda)$ is an eigenvector-value pair of $L^{\pi}$ if and only if $(Pv,\lambda)$ is an eigenvector-value pair of $L$.
This corollary tells us how the spectrum of the quotient graph finds its way into that of the full graph. We refer to these eigenvectors from the quotient graph as the structural eigenvectors. We borrow this naming convention from network science, the relation to which is discussed in section V.

### v2

@@ Abstract
Almost equitable partitions (AEPs) have been linked to cluster synchronization in oscillatory systems, highlighting the importance of structure in collective network dynamics. We provide a general spectral framework that formalizes this connection, showing how eigenvectors associated with AEPs span a subspace of the Laplacian spectrum that governs partition-induced synchronization behavior. This offers a principled reduction of network dynamics, allowing clustered states to be understood in terms of quotient graph projections. Our approach clarifies the conditions under which transient hierarchical clustering and multi-frequency synchronization emerge, and connects these dynamical phenomena directly to network symmetry and community structure. In doing so, we bridge a critical gap between static topology and dynamic behavior-namely, the lack of a spectral method for analyzing synchronization in networks that exhibit exact or approximate structural regularity. Perfect AEPs are rare in real-world networks since most have some degree of irregularity or noise. We define a relaxation of an AEP we call a quasi-equitable partition at level $\delta$ ($\delta$-QEP). $\delta$-QEPs can preserve many of the clustering-relevant properties of AEPs while tolerating structural imperfections and noise. This extension enables us to describe synchronization behavior in more realistic scenarios, where ideal symmetries are rarely present. Our findings have important implications for understanding synchronization patterns in real-world networks, from neural circuits to power grids.
The Kuramoto model is a network oscillator model useful for describing collective synchronization over a network of interacting entities, with applications from neuroscience to engineering to the social sciences. In this work we explore the relationship between the structure of the underlying network and the clustered synchronization that can occur in the dynamics. We look at the underlying network structure through the lens of Almost Equitable Partitions (AEPs), which are defined by regularity in their connectivity and relate to common notions of network community structure. The existence of this kind of partition is reflected in the eigenvectors of the network’s matrix representation as a graph Laplacian. We use this to relate the theory of community detection and symmetries of networks to the synchronization dynamics of the Kuramoto model by building a description in terms of these eigenvectors. When treated in this way, the question of collective synchronization becomes one of the population dynamics of eigenvectors. This relationship allows us to shed light on transient cluster synchronization, multi frequency clustering, and the conditions under which they can occur.
@@ I Introduction
The study of complex networks and their dynamical behaviors has been a fertile ground for understanding synchronization phenomena across different fields. From meso-scale neural activation, to power flow networks, numerous complex networks exhibit interacting oscillatory behavior. In the study of these systems, it is often natural to ask what sort of synchronization behaviors can occur. In recent years, there has been evidence from neuroscience of an unexplored relationship between resting state clusters in the brain and the Laplacian eigenvectors of functional connectivity networks.Atasoy, Donnelly, and Pearson (2016) While existing studies increasingly recognize the role of non-pairwise interactions in shaping complex neuronal synchrony,Majhi et al. (2025) they fall short of explaining the correspondence between functional clustering and structurally defined eigenvectors, particularly in the presence of high amounts of noise. Our work approaches this question from the perspective of network oscillatory models, closing this gap between structural and functional clusters.
Pioneering work by KuramotoKuramoto (1984) on what is now called the Kuramoto model, brought fundamental insights into how coupled oscillators spontaneously synchronize, and laid the groundwork for subsequent investigations into network-level synchronization dynamics.Strogatz (2000) The Kuramoto model and its derivatives are used as an abstraction of coupled periodic behavior for study. This model represents vertices as oscillators and couples their differential equations over the edges of the network. Consider a weighted graph $G$ with adjacency matrix $A$. Let $\sigma\in\mathbb{R}_{\geq 0}$ be a global coupling constant, and for every vertex $i$, define the natural frequency as $\omega_{i}\in\mathbb{R}$. Then the phase of vertex $i$ over time is determined by the differential equation
@@
The synchronization dynamics of coupled oscillator models can take many forms. In this paper we explore full synchronization and clustered dynamics in the context of the Kuramoto model, but in more general coupled oscillator models, there exist a broad range of hybrid synchronization patterns, as reviewed by Parastesh et al.,Parastesh et al. (2021) studying chimera states emerging from structurally heterogeneous or imperfectly symmetric networks. In parallel, studies have also demonstrated that cluster synchronization can arise not only from network symmetries but also from nonlinear coupling effects. For instance, Skardal et al.Skardal, Ott, and Restrepo (2011) and Gong and PikovskyGong and Pikovsky (2019) have shown that higher-order harmonic terms in the coupling function can induce multi-cluster dynamics even in fully connected networks without structural heterogeneity.
@@
Naturally, the synchronization characteristics of a coupled oscillator model, and the Kuramoto model specifically, are informed by the structure of the graph it is placed upon. However, the nature of this relationship is unclear, particularly due to the multitudinous structures a graph can take. Recent advances in the study of coupled oscillator dynamics have progressively refined our understanding of how graph structure influences synchronization potentialNguyên, Lee, and Stolz (2025); Schaub et al. (2016); Kato and Ishii (2023) and have proven that Almost Equitable Partitions (AEPs) facilitate cluster synchronization via their quotient graph. Specifically, the characterization of synchronized clusters in networked oscillators has highlighted the intricate relationship between graph structure and dynamical coherence.Abiad, Hojny, and Zeijlemaker (2022); O’Clery et al. (2013) Simultaneously, in more general signal processing literature, a novel approach has been emerging, which focuses on the discretization of the Fourier analysis framework to better study the fundamental tradeoff between a signal’s distribution across the vertices of a graph and in its spectral domain.Shuman et al. (2013)
@@
Despite these significant advances, a crucial gap persists between the spectral properties of AEPs, studied in graph theory and network science, and the cluster synchronization they allow in complex networks. Traditionally, their spectral properties have proven to be crucial for the detectability of AEPs (community structure) on graphs,Zhang and Sun (2021); Krzakala et al. (2013) but have till this point found little purchase in the space of signals constrained to the vertices of graphs and importantly for us, the dynamics of cluster synchronization. Foundational studies by Schaub et al.,Schaub et al. (2016) O’Clery et al.O’Clery et al. (2013) and Kato et al.Kato and Ishii (2023) have pointed to the importance of network structure and specifically Almost Equitable Partitions in the formation of clustered dynamics. However, these dynamical insights on the structurally equitable or near-equitable configurations have yet to be tied into the corresponding spectral theory of these structures for analysis.
@@
Building upon spectral graph theory and drawing from recent developments in network synchronization research, we propose a generalized framework that extends traditional equitable partition analysis. Researchers like Schaub et al., and Thibeault et al.Schaub, Li, and Peel (2023); Thibeault et al. (2020); Thibeault, Allard, and Desrosiers (2024) have provided foundational insights into matrix reduction and system dimensionality that inform our approach. At the same time, symmetry-based approachesCho, Nishikawa, and Motter (2017); Zhang and Motter (2020); Pecora et al. (2014); Khanra et al. (2022) have highlighted how both explicit and hidden symmetries in network topology can dictate the formation and stability of synchronized clusters.
However, a unifying spectral framework that links approximate structural patterns to emergent dynamical behavior has been missing. Our work addresses this gap by showing that eigenvectors associated with quotient graphs induced by almost equitable partitions (AEPs) form a spectral subspace that governs the coarse-grained dynamics of clustered synchronization. By leveraging eigenvector interpretation techniques,Shuman et al. (2013) we extend this framework to define quasi-equitable partitions at level $\delta$ ($\delta$-QEPs), which may retain dynamical relevance even when exact AEP structure is disrupted by noise or irregularity.
@@
In the sections that follow, we build the necessary foundation for an understanding of cluster synchronization in the Kuramoto model, entirely through the information encoded in its graph Laplacian. We begin by considering the structural information within the Laplacian’s spectrum, and particularly the role of partitions therein (Section 2). We then discuss the relationship between the spectrum of the Laplacian and its dynamics. Following this background, we present a novel spectrally motivated description of cluster synchronization, and discuss the stability of these properties with weighted graphs (Section 3). The presented spectral description of clustered dynamics allows for new and interesting applications. We explore the implications for multi-frequency clustering, and modeling with antisymmetric phase lag. Lastly, Section 5 offers concluding remarks and future research directions.
@@ II The Spectral description of Synchronized Clusters
The field of spectral graph theory involves itself with the linear algebra of graph representations and the relationship with graph structure. This crucial foundation for the field of network analysis allows us to leverage matrix algebra tools, built for computing, to understand graphs at scales we cannot feasibly study combinatorially.
Matrix representations of graphs come in many forms, and we present a few with relevance to our application. Let $G=(V,E,w)$ be a weighted connected graph with $n$ vertices $v_{1},\dots,v_{n}\in V$ and $m=|E|$ edges $E\subset\{\{u,v\}\mid u,v\in V\}$, where $w:E\to\mathbb{R}^{+}$ is the weight function for the edges. Perhaps the most natural description of this graph comes in the form of the adjacency matrix $A\in\mathbb{R}^{n\times n}$, whose $ij$th entry is $w(\{v_{i},v_{j}\})$ if $\{v_{i},v_{j}\}\in E$ and 0 otherwise.
Another important representation comes in the form of the signed incidence matrix, also known as the boundary matrix. This matrix ignores weight assignments, and requires an arbitrary orientation on the edges of our graph. For simplicity, we induce an orientation from the vertex order. Define the signed incidence matrix as $B\in\mathbb{R}^{n\times m}$ such that $b_{ie}=1$ if edge $e=\{i,j\}$ and $i<j$, and $b_{ie}=-1$ if $i>j$. Finally, $b_{ie}=0$ if edge $e$ is not incident to vertex $i$. This boundary matrix serves as a fundamental building block of the spectral form of the Kuramoto equation shown later. We will see that when viewed as an operator, it offers a bridge between vertex values and edge values.
Finally we introduce the matrix we pay particular attention to in this paper, the graph Laplacian. We define the graph Laplacian as L=DAL=D-A, where $A$ is the adjacency matrix and $D\in\mathbb{R}^{n\times n}$ is the diagonal matrix of the row sums of $A$. The entries of $D$ are the sums of incident edge weights. In the case of uniform weight $1$ (an ’unweighted’ graph), they equal the degrees of the corresponding vertices. The graph Laplacian is a cornerstone in spectral graph theory and has numerous applications in network analysis, clustering, and graph-based dynamical systems.Newman (2018)
@@
It is well-established that the eigenvectors of the Laplacian play a crucial role in revealing organizational structure in graphs, particularly in terms of their symmetry and connectivity patterns. For example, the multiplicity of the zeroth eigenvalue reveals the number of connected components and, in the case of multiplicity $1$, its corresponding normalized eigenvector is $v_{0}=\frac{1}{\sqrt{N}}\mathbb{1}$.
@@ II.1 Spectrum and Almost Equitable Partitions
One frame under which to understand symmetries of graphs or groupings of structurally similar vertices, is almost equitable partitions (AEP) of a graph, also known as externally equitable partitions.
We define an AEP on an unweighted graph before proceeding to our weighted case. Let $N(v)$ denote the set of neighbors of $v\in V$. An AEP is a partition of the vertex set $\pi=\{V_{1},\dots,V_{k}\}$ such that for $V_{i},V_{j}\in\pi$ where $i\neq j$, there exists $d_{ij}\in\mathbb{N}$ such that for all $v\in V_{i}$, we have that $|N(v)\cap V_{j}|=d_{ij}$ (See figure 1 B.1). In other words, the partition cell a vertex is in determines the number of neighbors that vertex has in each other partition cell. Such a partition encodes a grouping of vertices that play the same ‘role’ in the connectivity of the graph between the delineated partition cells. Note that the almost equitable partition ignores any connectivity within a partition cell, the case in which this too is regular is called an equitable partition (EP). The lack of constraints within partitions cells of an AEP allow for them to exist at multiple, overlapping, scales.
One can generalize this notion and define an AEP in the context of a weighted graph.Abiad, Hojny, and Zeijlemaker (2022) Let $G=(V,E,w)$ be a weighted graph with weight function $w:E\to\mathbb{R}^{+}$. Then a weighted almost equitable partition of $G$ is a vertex partition $\pi=\{V_{1},\dots,V_{k}\}$ such that for $V_{i},V_{j}\in\pi$ with $i\neq j$, there exists $d_{ij}\in\mathbb{R}^{+}$ such that for all $v\in V_{i}$ we have $\sum_{w\in N(v)\cap V_{j}}w(\{v,w\})=d_{ij}$. In other words, a weighted AEP is again a vertex partition with regular inter cell connectivity, except now the regularity comes in the form of a constant sum of edge weights rather than a constant number of neighbors. Note that we recover the original definition of an AEP with a weight $1$ uniform graph. For the rest of this paper our discussions on AEPs refer to the weighted AEP unless otherwise specified.
We can use this notion of vertex partitioning, or vertex role assignment, to coarse-grain the graph, summarizing the implied structure in what is known as the quotient graph (See figure 1 C.1). Consider a graph $G$ with arbitrary partitioning $\pi$. Define the indicator matrix of $\pi$ as $P^{\pi}\in\mathbb{R}^{n\times k}$ such that $P^{\pi}_{il}=1$ if vertex $i$ is in partition cell $V_{l}$, and 0 otherwise. Note we omit the $\pi$ superscript when the choice of partition is clear. The rows of $P$ correspond to the vertices of $G$ and its columns correspond to the partition cells of $\pi$. Now given an $n$ by $n$ matrix representation of a graph like the graph Laplacian (or adjacency), $L$, we define the quotient Laplacian (or adjacency) matrix of $L$ with respect to the partition $\pi$ as
@@
Figure 1: In A.1, we show the original graph, and in A.2, its associated graph Laplacian matrix. The graph admits an almost equitable partition (AEP) into three cells, visually suggested in B.1 by vertex coloring. In C.1, we display the quotient graph induced by the AEP, and in C.2, the eigenvectors of its quotient Laplacian. The structure of this quotient graph is reflected in eigenspectrum of the graph Laplacian B.2, with the highlighted eigenvector revealing the partition through its block structure. In particular, this structural eigenvector is the scaled counterpart of the highlighted eigenvector in C.2. This highlights the alignment between the structural features of the full graph and its coarse representation.
The quotient graph provides a powerful tool for coarsely analyzing the structure of the original graph, while preserving connectivity information relevant to the partition. Indeed, cluster synchronization can be understood as a reduction of the dynamics onto such a quotient graph.Schaub et al. (2016) The preservation of information is made clear in the analysis of the quotient Laplacian in the context of an almost equitable partition. Notice, in the case of an almost equitable partition $\pi$, the off diagonal blocks of the induced block partition of $L$ have average row sums exactly equal to one another. This implies that the off diagonal entries of $L^{\pi}$, which correspond to the average of these row sums, preserve all the information of inter cell connectivity. The following theorem states this algebraically.
@@
Brouwer and Haemers (2012); Cardoso, Delorme, and Rama (2007); Aguilar and Gharesifard (2016) Let $L$ be the Laplacian of a graph $G=(V,E,w)$, and $\pi=\{C_{1},\cdots,C_{k}\}$ be a partition of $V$.
The partition $\pi$ is an almost equitable partition if and only if the image of $P^{\pi}$ is invariant under $L$, which means there exists a matrix M such that $LP=PM$. Moreover, the solution to this Sylvester matrix equation is unique, namely $M=L^{\pi}=(P^{T}P)^{-1}P^{T}LP$.
Theorem 1 comes from a more general theory of the tight interlacing the eigenvalues of $L^{\pi}$ have with respect to the eigenvalues of $L$.Brouwer and Haemers (2012) It is helpful to note that the expression for the quotient Laplacian $L^{\pi}$ can be written using the Moore-Penrose pseudoinverse as $L^{\pi}=P^{+}LP$. This follows from Penrose’s 1955 classical paperPenrose (1955) and offers an alternative derivation of Theorem 1. Then given, $(\lambda,v)$, an eigen-pair of $L^{\pi}$, it follows
@@
Brouwer and Haemers (2012); Aguilar and Gharesifard (2016) Let $\pi$ be an almost equitable partition of a graph, and $P$ the indicator matrix associated to $\pi$. Then $(v,\lambda)$ is an eigenvector-value pair of $L^{\pi}$ if and only if $(Pv,\lambda)$ is an eigenvector-value pair of $L$.
This corollary tells us how the spectrum of the quotient graph finds its way into that of the full graph. This relationship is exemplified in figure 1. The same is true for all AEPs and their corresponding quotient graphs, admitting any structure within partition cells and not just the fully connected one shown in this toy example. We refer to these eigenvectors from the quotient graph as the structural eigenvectors. We borrow this naming convention from network science, the relation to which is discussed in section V.